OpenStudy (anonymous):

How do you do this linear equation with three variables?: {-2x+3y-5z=3 5x-7y+12z=-8 X-y+2z=-2

OpenStudy (anonymous):

Find the value of one variable x, y or z from any one equation and then put it in other two equations, you will get two equations containing only two variables, then solve this further to find the two variables first..

OpenStudy (anonymous):

Suppose we find y from Third equation: \[x - y + 2z = -2 \implies y = x + 2z + 2\] Now put this value of y in other two equation one by one..

OpenStudy (anonymous):

\[-2x+3y-5z=3 \implies -2x + 3(x + 2z + 2) - 5z = 3\] \[-2x + 3x + 6z + 6 = 3 \implies x + 6z = -3\]

OpenStudy (anonymous):

Similarly in second one: \[ 5x-7y+12z=-8 \implies 5x - 7(x + 2z + 2) + 12z = -8\] \[5x - 7x - 14z - 14= -8 \implies -2x - 14z = 6 \implies x + 7z = -3\]

OpenStudy (anonymous):

Now you have got two equations: \[x + 6z = -3\] \[x + 7z = -3\] Solve to find x and z..

OpenStudy (anonymous):

Can you do this @K.marie ??

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

Solve it and post your answer so that I can check..

OpenStudy (anonymous):

X=-39 y=-49 z=-6

OpenStudy (anonymous):

How?? No.. Just solve those two equations..

OpenStudy (anonymous):

Oh then z=-6 and when I plug that back in I get x=-39

OpenStudy (anonymous):

How z = -6 ??

OpenStudy (anonymous):

Multiply x+6z=-3 by -1 and add to x+7z=-3 to get z=-6

OpenStudy (anonymous):

By multiplying you will get: -x - 6x = 3 x + 7z = -3 Now add them what you got?

OpenStudy (anonymous):

X gets cancelled and the 6 is actually a z so that becomes one z

OpenStudy (anonymous):

Yes that becomes one z.. So what is -3 + 3 = ??

OpenStudy (anonymous):

ON RHS what will be the result??

OpenStudy (anonymous):

0 so it's a no solution

OpenStudy (anonymous):

Hey, it is having solution..

OpenStudy (anonymous):

But here z = 0, now find x and y..

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

X=-3 and y=-1

OpenStudy (anonymous):

Yep, now you are right.. So, x = -3, y = -1 and z = 0..